UT September 06, 2019 - Restart Night #5


OSA: S. Allanson

ISp: E. Solheid

ISA: B. Rothberg, O. Kuhn (Tucson)

Mtn Manager: P. Hartley

Technical Expert: J. Hill (Tucson)

Plan for Tonight:
  • LUCI - AGW2 has been repaired
    • AGW checkout and alignment
      • adjust M3 to minimize pupil wobble (critical) - CANNOT DO BECAUSE OF IDL ISSUE
    • AGW-1 and AGW-2 Transform - high (2x2hrs - requires initial data, then followup to confirm transform) - CANNOT DO BECAUSE OF IDL ISSUE
      • LUCI1/AGW1 rotator ZP looks good - no need to update PCS config
      • LUCI2/AGW2 rotator ZP looks off by ~0.2 degrees (see previous nights log)
    • Imaging check - dither outside of field
    • non-sidereal check imaging - verification of 2019A TCS/PCS/GCS upgrades
    • Narrow band twilight flats (low priority)
  • LBCs (contingent on LBCB, new LBCR striping, clear and/or good seeing)
    • Super-focus sequenece PA=0, 180 (1hr)
    • Test Science OB (sidereal)
    • Standard photometric fields
    • Spectro-photometric standards
    • Twilight Flats
  • PEPSIPFU wake-up checks (high)
  • DONE M3 alignment for PFU pupil (1 hr)
  • MODS
    • check MODS1 probe w/rt rotator center
    • NS LS tests
    • DONE Imaging test OB
    • field aberrations (align M2 with M1)(MODS or LUCI)
    • DONE twilight flats


UT 01:40 - initi all on LUCIs, switching to MODS at start because we cannot do AGW transform or M3 alignment. IDL scripts are not working at the start of the night. (Doug fixed the problem with IDL later.)

UT 01:41 - LBCs up, MODS starting up

Open

UT 02:00 - starting MODS skyflats (had a co-pointing error from TCS/PCS and had to reauthorize)

UT 02:20 - finished skyflats

UT 02:25 - pointing and collimating MODS-Bino, then will do dither test with MODS imaging

UT 02:31 - leftover astigmatism on DX (~7000) which should not have been there, had to be cleared out - turns out to be leftover from AdSec flat issues in the spring.
(JMH corrected the M1 Z6 LUT back to that corresponding to a zenith flat, and added a note to IT 7629.)

UT 02:37 - Stone O with lots of dithers

UT 02:44 - MODS2R IMCS lock timed out

UT 02:51 - MODS2R timeout

UT 03:26 - MODS1R is 12 exposures behind on readout - will run red fitsflush when next readout is done

UT 03:40 - now testing MODS syncoffsets - 4pt dither - same field

UT 03:45 - MODS2R timeout on imcslock. Did an "r" to retry, and that was successful.

-but left side timed out on syncoffset after 60 seconds (so it said, but the IMCS timeout takes longer than 3 minutes, and it is possible that just the print statement is incorrect).

-proceeding OK now, but O. Kuhn checking software regarding whether syncoffset timeouts are the same as preset timeouts (450 seconds) or whether the print statement is incorrect. NB: reading through the execMODS script, presetTO is defined to be 300 sec, but for syncpreset, the timeout = 1.5*presetTO (this was added as documented in IT 7774) while for syncoffset, timeout = presetTO.

UT 04:01 - MODS1 B and R 0039 images are trailed - it looks to be behind and/or exposed on an offset. The whole while this script was running, the MODS1 commands and GUI were behaving oddly. MODS1/2 GUIs showed different x,y offsets and the MODS1 GUI "send offset" was orange, as though the offset was queued up and ready to send once the MODS2 script had reached that point. In the MODS1 script running window, the commands/feedback were not exactly as in the MODS2 window. After seeing the trailed pair of MODS1 images, decided to abort. This reinforced the lesson: Use "i" to ignore a script error at your own risk. Rick said as much in IT 7774.

-going to cancel and restart the script

UT 04:04 - resending script

-MODS2 timed out on object name

-Timed out again on name (both MODS)

UT 04:12 - resending preset - the problem was the IIFGUI was "backed up" with multiple requests - it was hung waiting for a sync offset on one side and a sync preset on the other (CTRLC stopped the sync scripts for Bino-MODS, but we didn't pass on clearly the request to cancel the preset/offsets before sending a new synchronous preset).

-red imcslock failed on right side, sync offset waiting on left

-still timing out after 60 seconds? or is message that pops up in terminal incorrect - when hitting Ignore the script does not handle it correctly and it leads to trailed images or other issues. (Note added to IT 6939)

-going to run script from the execution block - after the preset in the imaging script - and after editing it to comment out the WFS wait:

UT 04:30 -CTRL C because forgot to send it as binocular - redoing that now.
execMODS --mods1 --bino -e edited_script.txt
execMODS --mods2 --bino -e edited_script.txt

UT 04:48 - finished with syncoffset test. Script finally ran through without problem - going to do MODS NS test.


NS Observation with MODS

NS target: 2014 AF51 - potential hazardous asteroid - LBT safety patrol

-first pair of exposures were done without WFS

-target is right on quadrant boundary on MODS2B, OK in MODS2R, on four quad boundary on MODS1B and MODS1R

-sending an offset on each side

-2nd set of exposures target is in a clean area away from bad columns or quad boundaries

UT 05:24 - finished NS testing


Field Aberrations w/MODS (attempted)

-postponed due to clouds


LUCI Dither test Imaging

UT 05:32 - switching to LUCIs

UT 05:42 - moving to M52 dither test

UT 05:51 - first need to verify AGW2 WFS is working correctly

-collimated on 7th mag star

UT 05:55 - sending preset to M52 field - bino dithers 300" off field and cycles through KJH

-R=14 too faint, trying R=12.3

L1 0004-

L2 0004-

UT 06:06 - took 2 images, lost guiding on 3rd exposure - put one blind in FW1 and will just wait

-at the same time the faintness was useful for checking the alignment of the star in the WFS pupil

UT 06:11 - switching to PEPSI now - blinds in LUCIs and flexure off


PEPSIPFU wfs pupil alignment check

UT 06:15 Moving to PEPSIPFUs in fairly thick clouds. Steve had to reboot agwcam 7 and 8 and restart both GCSs.

DX needed a bunch of manual Z7 coma correction (need to check that later).

UT 06:38-06:45 Converged pupils on wfs at this time without any change to M3 position or GCS parameters.


LUCI - attempting dither test again

UT 06:53 - collimated - attempting with R=11.9 guidestar and less on dithers

L1 0006

L2 0006

-made the offset and came back but guiderstar is super faint

UT 07:02 - lost the guidestars again

UT 07:05 - Switching to NGC 7585 - RA 23 Dec -04 - looks to be in the hole

-dither test +/-30" in field +/-300" off field dither

-seems to be working, but top of WFS on SX is cutoff (guidstar is 4.5' away), left side of WFS on DX is cut off (tertiary centered?)

-with clouds WFS showing ~0.5-0.6"

UT 07:47 - looks like we are staring into a cloud

UT 07:52 - Finished NGC 7585 - K,H,J 5 on 4 off x2 LUCIs (60sec per)

-guidestar fading at end of last exposure


M3 Alignment at LFBG/RFBG with Pupil Wobble measurements

UT 07:54 on-axis ACTIVE preset in clouds

Procedure lives in ComTools

working in /home/lbto/data/20190906_lfbg and 20190906_rfbg

SX M3 is at Abs Tip=-580 Tilt=-160 RZ=29.799
DX M3 is at Abs Tip=-664 Tilt=+80 RZ=-25.721

Pointing model change - IE/CA: SX => -56.0/-35.0; DX => -83.0/7.0

UT 07:58 pupil_collect, but forgot to stop wfsing the first time.

UT 08:01 pupil_collect

wobble
LFBG
.
left_wfsc000001_dd.fits
left_wfsc000002_dd.fits
left_wfsc000003_dd.fits
left_wfsc000004_dd.fits
# Shack-Hartmann Pupil Wobble data
# x_center   y_center   rotangle    imagename
#----------------------------------------------
    386.09    242.48    304.85   left_wfsc000001_dd.fits
    381.77    237.55    245.85   left_wfsc000002_dd.fits
    377.81    240.94    546.41   left_wfsc000003_dd.fits
    376.71    247.17    486.93   left_wfsc000004_dd.fits

$ xmgrace -par ../pupil_wobble.par -block pupil_07:59:53.wobble -type xyz -bxy 1:2:3 &

wobble
LFBG  (Error in log file -should be RFBG)
.
right_wfsc000001_dd.fits
right_wfsc000002_dd.fits
right_wfsc000003_dd.fits
right_wfsc000004_dd.fits
# Shack-Hartmann Pupil Wobble data
# x_center   y_center   rotangle    imagename
#----------------------------------------------
    380.28    264.69    -59.32   right_wfsc000001_dd.fits
    376.85    265.25   -118.35   right_wfsc000002_dd.fits
    376.89    269.12    182.17   right_wfsc000003_dd.fits
    376.95    271.62    122.73   right_wfsc000004_dd.fits


UT 08:09 Had to close for clouds for a bit - which gave JMH and OPK time to learn to make proper plots of pupil wobble.


$ xmgrace -par ../pupil_wobble.par -block pupil_08:01:55.wobble -type xyz -bxy 1:2:3 &

DX Pupil Wobble Stats (post facto analysis on 20200903) - puzzling because these notes suggest that M3 didn't move between these 2 measurements
filename              Xavg    Yavg    Radius      pts
pupil_08:50:30.wobble 376.66 266.9525 9.3360524313 4
pupil_08:01:55.wobble 377.7425 267.67 6.40562838448 4

  • DX M3 Pupil Wobble Plot:
    Screenshot at 2020-09-03 21-28-04.png

UT 08:50 pupil_collect 08:50:24
again with fewer clouds. We haven't moved M3 yet. These results were consistent with the previous ones through clouds.

UT 09:05 SX dTip +100 (Abs -480)

SX pupil_collect 09:06:33

UT 09:15 SX dTip remains 100 dTilt +100 (Abs -60)

SX pupil_collect 09:16:53

UT 09:2x SX dTip +100 to 200 (Abs -380) Tilt back to 50 (-110 abs)

SX pupil_collect 09:28:11

UT 09:40 SX Tip to 300 (Abs -280) Tilt to 0 (Abs -160)

SX pupil_collect 09:41:37

UT 09:52 SX Tip to 400 (Abs -180) Tilt to -50 (Abs -210) THESE ARE THE NEW ADOPTED VALUES. With IE=-40 CA=-4

SX pupil_collect 09:52:58

UT 10:03 IE/CA: SX => -40.0/-4.0

xmgrace -par ../pupil_wobble.par -block pupil_08:50:24.wobble -type xyz -bxy 1:2:3 -block pupil_09:52:58.wobble -type xyz -bxy 1:2:3 -block pupil_09:28:11.wobble -type xyz -bxy 1:2:3 &

SX Pupil Wobble Stats (post facto analysis on 20200903)
filename              Xavg    Yavg    Radius      pts
pupil_09:52:58.wobble 383.0975 244.1 1.125022222 4
pupil_09:41:37.wobble 383.19 244.0675 2.8628787959 4
pupil_09:28:11.wobble 383.5325 243.175 5.07188081485 4
pupil_09:16:53.wobble 383.03 242.215 7.46744266801 4
pupil_09:06:33.wobble 382.505 241.705 7.39763475714 4
pupil_08:50:24.wobble 382.545 241.6425 10.4297926633 4
pupil_07:59:53.wobble 380.595 242.035 10.1146230775 4
  • SX M3 Pupil Wobble Plot:
    Screenshot at 2020-09-03 21-04-04.png

(JMH installed these absolute values for SX M3 into positions.conf on the evening of 20190907 UT.)
oss.side[0].terc.focal_stations.bentGregorianFront              map  0:-180.0   1:-210.0  2:0.0  3:+29.800
oss.side[0].terc.focal_stations.bentGregorianArgos              map  0:-180.0   1:-210.0  2:0.0  3:+29.807

Field Aberration Data for LFBG / RFBG

working in /home/lbto/data/20190906_lfbg and 20190906_rfbg

BS9188_field.log

10:07 on-axis ACTIVE preset to BS9188 - seeing is 0.6-ish

weird error where SX WFS blocked (see Night Log for the tcslog timestamp) - Steve resent preset

10:12 field_collect running on both sides in separate windows

We have NOT switched off DX high order zernikes, but it doesn't seem to be sending more (because GCS is taking the wfs images).

We are collimating in 1 iteration. Should I have tightened the tolerance from the default (300)? Of course, the seeing got worse after I wrote that.

10:30 crazy guide oscillation on DX, field_collect errors out
(See IT 7844 created tomorrow night.)

10:34 on DX restart: field_collect, start=4

10:59 end of field data collection

DX #| GStar | wfsc # | 201996 BS9188_field wfsc GCS
0 416
11 417-419
0 420-422
12 423-425
0 426-427
13 428-430
0 431

#| GStar | wfsc # | 201996 BS9188_field wfsc GCS
0 442-443
18 444-446
0 447-448
17 449-451
0 452-453
4 454-456
0 457-458
9 459-461
0 462-464
7 465-467

SX

#| GStar | wfsc # | 201996 BS9188_field wfsc GCS
0 401
11 402-404
0 405-406
12 407-409
0 410
13 411-413
0 414
18 415-417
0 418-419
17 420-422
0 423-425
4 426-428
0 429
9 430-432
0 433
7 434-436
0 437-439

#| GStar | wfsc # | 201996 BS9188_field wfsc GCS
0 444
11 445-447

SX Field Data Processing (by JMH on 20190907)

cd /home/lbto/data/20190906_lfbg/

edit BS9188_field.log to correct path and remove superfluous headers

iif_register, LFBG

field_make_files (ran OK)

field_process_set (processed all the wfsc images, but then failed with )
% Compiled module: FIELD_DETERMINE_FIELD.
% Compiled module: FIELD_DATA__DEFINE.
% Program caused arithmetic error: Floating illegal operand
% Illegal subscript range: SET.
% Execution halted at: FIELD_DETERMINE_FIELD  150 /home/lbto/idl/wfsc/field/field_determine_field.pro
%                      FIELD_PROCESS_SET   76 /home/lbto/idl/wfsc/field/field_process_set.pro
%                      $MAIN$          

JMH also has doubts about whether pupil center is sufficiently well specified. This data set should give a better center than what I originally measured.

JMH also wonders if it depends on NOT having two on-axis wfs images in a row (at different PA's).

DX Field Data Processing (by JMH on 20190907)

cd /home/lbto/data/20190906_rfbg/

edit BS9188_field.log to correct path and remove superfluous headers

iif_register, RFBG

field_make_files (ran OK)

field_process_set (got same error as SX side did)
% Compiled module: FIELD_DETERMINE_FIELD.
% Compiled module: FIELD_DATA__DEFINE.
% Illegal subscript range: SET.
% Execution halted at: FIELD_DETERMINE_FIELD  150 /home/lbto/idl/wfsc/field/field_determine_field.pro
%                      FIELD_PROCESS_SET   76 /home/lbto/idl/wfsc/field/field_process_set.pro
%                      $MAIN$          


LBC for Non-sidereal observation

UT - 11:00 Reconfiguring to LBC/LBC

-proceeding with NS observations of 2018YH for C. Veillet

-need to collimate and co-point first

UT 11:19 - slewing - however we got the co-pointing error, re-authorizing and re-sending

UT 11:23 - running dohybrid

UT 11:32 - clouds rolled in - can't see any stars in field

UT 11:37 - closing up - too many clouds and too close to sunrise - also clouds are super thick

LUCI JHK Bino Imaging flats

  • L1/L2 0037-0041 J lights off - agw on-axis don't use
  • L1/L2 0042-0046 J lights on - agw on-axis don't use

  • L1/2 0047-0051 J lights off
  • L1/2 0052-0056 J lights on
  • L1/2 0057-0061 H lights off
  • L1/L2 0062-0066 H lights on
  • L1/L2 0067-0071 K lights off
  • L1/L2 0072-0076 K lights on
Also taking MODS biases

UT 12:02 - noticed LBC-Red biases have striping again, 2nd morning in a row we see that

UT 12:07 - LUCI-2 put in safe config

UT 12:08 - LUCI-1 put in safe config

UT 12:11 - no LUCI data after #0039 is appearing in /Repository


-- BarryRothberg - 05 Sep 2019
Topic attachments
I Attachment Action Size Date WhoSorted ascending Comment
Screenshot at 2020-09-03 21-04-04.pngpng Screenshot at 2020-09-03 21-04-04.png manage 49 K 03 Sep 2020 - 21:33 JohnHill SX M3 Pupil Wobble Plot
Screenshot at 2020-09-03 21-28-04.pngpng Screenshot at 2020-09-03 21-28-04.png manage 42 K 03 Sep 2020 - 21:34 JohnHill DX M3 Pupil Wobble Plot
Topic revision: r16 - 03 Sep 2020, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback